40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8566     Entry time: Mon May 13 23:05:26 2013
Author: Koji 
Type: Configuration 
Category: LSC 
Subject: PRMI locking 

- Disabled MCL path in mcdown/mcupscript.

Nominal gain in mcdown/mcup was -50 and -100 respectively.

- Confirmed the stable lock was just because of the quiet seismic of the Friday night.

- Improvement of the PRM ASC servo
RG3.2 (3.2Hz Q=2 Height 30dB)
=>
RG3.2 (3.2Hz Q=10 Height 30dB) +  zero[f, 1, .5] pole[f, 2, 3] zero[f, 4.5, .5] pole[f, 3.5, 3]

Filter shape comparison is found in the second plot attached.

The resulting spectra (freerun vs controlled) is found in the first plot.

Nominal PRM ASC gain is +70

- Openloop TF measurement

OLTF PRCL 250Hz 30deg / MICH 200Hz 45deg

- REFL55/REFL33 phase adjustment (in lock)

REFL55 phase fine tune (95.25deg) (x1,x0.3)
REFL33 phase (-13.0deg) (x1, x2)

Attachment 1: 130513_PRC_ASC.pdf  58 kB  | Hide | Hide all | Show all
130513_PRC_ASC.pdf
Attachment 2: 130513_PRC_ASC_servo.pdf  38 kB  Uploaded Tue May 14 00:36:40 2013  | Show | Hide all | Show all
ELOG V3.1.3-